Is an event with a probability of 1 more or less likely than an event with a probability of 2?
It is more likely because it can exist. An event with a probability of 2 cannot exist.
How do you calculate mean and Median smaller then Standard deviation?
In the same way that you calculate mean and median that are greater than the standard deviation!
How do you draw a club from a deck of cards?
You have pimples? Me too! When i was in middle school i had a really clear face and once i started high school i began to get some pimples and now i have more especially on my forehead and chin, but they're small and there are scars now because i used to pop them and i don't know how to get rid of those little scars, and i found out that its mostly because i touch my face because i find them so annoying. what do i do to get rid of previous scars and the ones i have now?help in need of clear skin again.
What is the probability of rolling 6 and four in backgammon three times in a row?
The answer depends on how many times in total the dice are rolled. As the total number of rolls increases, the probability rolling a 6 and 4 three times in a row increases towards 1.
What is the probability of getting 4 ace?
To get four Aces it is necessary to draw several cards. There is no information on how many cards are drawn. Whether or not they are drawn randomly.
Give the reasons why sample are prefered to entire population?
When a population is too large, then a sample of the population would be taken to determine the information wanted. For example if you wanted to find out the average IQ of Americans you would not want to give every person living in the United States an IQ test, because that would be highly unrealistic, instead you would want to sample a random group of Americans from multiple backgrounds to get an average IQ of Americans.
The probability is 0.664
How common are fraternal twins?
In the US, the incidence of fraternal (dyzygotic) twins is 32 in 1000 live births.
What is the probability that flipping two coin will show two tails?
1 in two but they say the side with heads is slightly Heavier.
Is it acceptable if standard deviation is 8?
It may or may not be acceptable. If the mean is 12, then no it is not acceptable. If the mean is 1000, then it may be acceptable depending on the criteria given.
If the probability that an event will occur is 1 then the event is certain to occur?
Yes, the probabilty scale is from one to zero so is the probability is 0 it will definitely not occur
Does smoking weed produce more girl babies than boys?
There is no evidence in support of the assertion.
What is the probability of getting a red queen and a 4 on the die?
Pr(Red queen) = 2/52 = 1/26 = 3.85%
Pr(4 on a die) = 1/6 = 16.67%
Pr(Red queen AND 4 on a die) = 1/26 * 1/6 = 1/156 = 0.64%
What if your great grandma had twins and your dad is a twin will you have twins?
Most likely not, as it shows to skip a generation. It skipped your grandma, so by logic, the pattern is yes-no-yes-no. You would the the fourth no.
It will also depend on which side of the family the twins were on. The propensity to have fraternal twins can be inherited as a female trait that causes more than one egg to be dropped at a time.
Identical twins have no genetic relationship that has been found to date.
The probability that all four of the cards are different suits is 13/17 * 13/25 * 13/49 = 2197/20,825 or 10.54982%.
Where can one purchase a rolling utility cart?
There are a few places that offer a variety rolling utility cart for sale. The cheapest place with the largest collection is the online marketplace Amazon.
How probabilities are used in business?
Well, it just does. If you have a problem with that, then that's your problem, don't kill wikianswers.
What is the probability of randomly drawing a face and heart from a standard deck of cards?
There are 52 cards in a deck. 1/4 are hearts. 3 are heart faces. So the probability of drawing one card that is a heart face is 3/52.
The chance of drawing a face card as one card, and a heart as another is
1/4(heart)*3/13(face)=3/52 as well
What is selective distribution?
Selective distribution is a type of retail strategy. It involves making specific items for sale only to specific geographical markets.
One must first consider the number of ordered ways that 4 students may be chosen, and then divide that by the number of ways the four may have been ordered to create a number of distinct groups.
For the orders, there are 30 choices for the first, 29 for the second, 28 for the third, and 27 for the fourth. However, there are a 4 ways within the four to pick the first, 3 ways for the second, 2 for the third, and 1 left for the fourth. The answer may be expressed by the following:
30 x 29 x 28 x 27
______________
1 x 2 x 3 x 4
which evaluates to 27,405 committees.
How do you solve a standard normal distribution?
You do not solve a standard normal distribution. It is not a question nor an equation or inequality to be solved. You can answer questions using the standard normal distribution but what you do depends on the question and on what information is given.
How many queens of diamonds in a deck of cards?
In an ordinary 52-card deck, there is one Queen of Diamonds.